You are considering an investment in Keller Corp's stock, which is expected to pay a dividend of $1.75 a share at the end of the year (D1 = $1.75) has a beta of 0.9. The risk-free rate is 3.1%, and the market risk premium is 5.5%. Keller currently sells for $42.00 a share, and its dividend is expected to grow at some constant rate g. Assuming the market is in equilibrium, what does the market believe will be the stock price at the end of 3 years? (That is, what is ?) Round your answer to two decimal places.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=3.1+(5.5*0.9)=8.05%

Required return=(D1/Current price)+Growth rate

0.0805=(1.75/42)+Growth rate

Growth rate=0.0805-(1.75/42)

=0.0388333333

P3=Current price*(1+Growth rate)^3

=$42*(1+0.0388333333)^3

=$47.09(Approx).